OpenStudy (anonymous):

Find the value of x if the area of the trapezoid is 54.

OpenStudy (callisto):

excuse me, but is there a diagram or some descriptions?

OpenStudy (anonymous):

yes sure one sec

OpenStudy (anonymous):

do you know the formula for the area of a trapezoid?

OpenStudy (anonymous):

Area = (1/2)(base1 + base2)*(height)

OpenStudy (anonymous):

just plug everything we know into the formula and solve for the unknown (x)

OpenStudy (anonymous):

54 = (1/2)(7 + x)(6) can you take it from here? solve for x....

OpenStudy (anonymous):

so would be it 1/2(x+7)6

OpenStudy (anonymous):

don't forget about the 54... otherwise you won't have an equation to solve.... :)

OpenStudy (anonymous):

i got an answer... you want a verification....

OpenStudy (anonymous):

yes verification

OpenStudy (anonymous):

and could you help me with this question : Determine the area of the figure

OpenStudy (callisto):

Area of triangle = 1/2 (base)(height) = 1/2 (4)(20) = ... Area of trapezium = 1/2 (upper base + lower base)(height) = 1/2 (12+20) (3) = ... Area of the figure = Area of triangle + Area of trapezium = ?

OpenStudy (anonymous):

40 and?

OpenStudy (callisto):

1/2 (12+20) (3) = 1/2 ( 32)(3) =?

OpenStudy (anonymous):

196?

OpenStudy (callisto):

Nope :|

OpenStudy (anonymous):

booo what is it?

OpenStudy (anonymous):

136?

OpenStudy (callisto):

Area of the figure = Area of triangle + Area of trapezium = = 1/2 (4)(20) + 1/2 (12+20) (3) = ?

OpenStudy (anonymous):

88, 100, or 70, or 136?

OpenStudy (anonymous):

ITS 88!

OpenStudy (anonymous):

Okay last question : For a 270° counterclockwise about the origin, multiply the vertex matrix on the left by

OpenStudy (callisto):

The question is not complete :|